# Interactive Perception for Deformable Object Manipulation

### v1

@@ I Introduction
Figure 1: An example of Interative Perception. The perceiver (camera) is moved to a new viewpoint while the actor (end-effector) opens the bag for better perception of in-bag object.
To this end, we propose to address interactive perception for deformable objects in a partially observable Markov Decision Process (POMDP) framework. We focus on bag-like objects and a dual-arm agent setup with an active camera similar to Fig. 1. Our aim is to study the feasibility of solving DOM for IP under the POMDP formulation and whether active vision can facilitate the solution. To leverage the regularity in coupling the perceiver and the actor, our approach proposes to dynamically construct a motion subspace, Dynamic Active Vision Space (DAVS), based on object features for an efficient search. Specifically, we introduce a manifold-with-boundary formulation to characterize a compact camera action subspace to constrain the potential next view. We provide detailed algorithms to calculate the manifold and integrate it into a general action search framework such as reinforcement learning (RL) in solving the POMDP. Our results in simulation with a bag object with similar but unseen dynamics highlight the importance of active vision capability from the controlled camera and the superior performance due to DAVS compared to standard baseline methods. We show the proposed method is robust for a successful transfer to the real hardware with a moderate fine-tuning on a dual-arm setup.
@@
  • We conduct extensive studies in both simulation and real-world experiments to demonstrate the effectiveness of our methods in challenging DOM scenarios, in particular, we show that our method generalizes well to objects with different dynamical properties.
@@ II Related Work
Figure 2: [Left] Illustration of the proposed framework — a subspace of camera action is constructed and represented with manifold with boundary, accounting for the coupling with end-effector motion via the structure of interest (SOI) on deformable objects. [Right] SOI on deformable objects for different tasks (e.g., opening, hanging, dressing).
@@ III-A Action Space Factorization and Manifold with Boundary
Considering an IP setup exemplified in Fig. 1, we exploit coupled camera motion $a^{cam}$ and end-effector motion $a^{e}$ to factorize the entire action distribution conditioned on observation $o$ as $\pi(a^{cam},a^{e}|o)=p(a^{cam}|a^{e},o)p(a^{e}|o)$. This factorization implies a possibility of sampling $a^{cam}$ from a subspace depending on $a^{e}$ and hence an efficient exploration of the original action space. We use manifold with boundary ${M}$ [34] as a tool to describe this subspace and its construction. The introduction of ${M}$ essentially imposes constraints, and thus structure/regularity, on the original camera action space. To this end, we augment the POMDP $({S},{O},{Z},{T},{R},\gamma)$, such that $z\in{Z}={M}\times{A}^{e}$ is the constrained action space subjecting to ${M}$. The new constrained IP problem can be written as
@@
III-B Dynamic Manifold with Boundary-based Active Vision Space
@@
Figure 3: Illustration of the process of dynamic active vision space (DAVS) generation. [Left] We extract the SOI points at each time step $t$ and convert them to projected 3D SOI (blue) points through ray tracing, on the original camera action manifold. [Right] We compute the manifold with boundary based on projected SOI and current camera position by Algorithm 2
Figure 3: Illustration of the process of dynamic active vision space (DAVS) generation. [Left] We extract the SOI points at each time step $t$ and convert them to projected 3D SOI (blue) points through ray tracing, on the original camera action manifold. [Right] We compute the manifold with boundary based on projected SOI and current camera position by Algorithm 2
@@ 3       Initial parameter setting for active vision ;
5             SOI Modeling $\mathbf{X}(t)\leftarrow f_{SOI}(\boldsymbol{o}_{k})$ ;
6             Sample policy end-effector action $\boldsymbol{\alpha}_{k}^{e}\sim\pi(\cdot\mid\boldsymbol{s}_{k})$ ;
@@
9             Sample policy camera action $\boldsymbol{\alpha}_{k}^{cam}\sim\pi(\cdot\mid\boldsymbol{s}_{k},\boldsymbol{v}_{1},\boldsymbol{v}_{2})$ ;
@@ 10             Apply the control camera action $\boldsymbol{\alpha}_{k}^{cam}$ to the environment ;
12             Observe the next state $\boldsymbol{s}_{k+1}$ and reward $r_{k}$ from the environment ;
13             Provide the transition tuple $(\boldsymbol{s}_{k},\boldsymbol{\alpha}_{k},\boldsymbol{s}_{k+1},r_{k})$ to the RL algorithm ;
@@ 2 Generate refined active vision space, ${M}^{\prime}(t)$ and its boundary $\partial{M}^{\prime}(t)$ ;
In addition to the centroid, we find two points $P_{1}$ and $P_{2}$ on $\partial{M}^{\prime}(t)$, such that the geodesic path passing through them is locally perpendicular to the one passing through $P_{0}$ and $O_{0}$. As a result, the DAVS is obtained by enclosing the three paths, $\gamma_{\overset{\frown}{P_{0}P_{1}}},\gamma_{\overset{\frown}{P_{1}P_{2}}},\text{and}\gamma_{\overset{\frown}{P_{2}P_{0}}}$. The full process is summarized as Algorithm 2.
@@ III-C Exploiting DAVS in Action Exploration
Figure 4: Quantitative evaluation on CubeBagClean. The first row represents the episode length ($\downarrow$) for solving 4 tasks across different combinations of random (Rand) and fixed (Fix) camera (Cam) and end-effector (EE) birth modes. The second row shows the total discounted rewards ($\uparrow$).
Figure 5: Quantitative evaluation on CubeBagObst subtasks. The first two figures represent the episode length ($\downarrow$) and reward ($\uparrow$) for the scenario with fixed camera (Fix Cam) and random end-effector (Rand EE) birth modes. The third and fourth figures are episode length ($\downarrow$) and reward ($\uparrow$) with random camera (Rand Cam) and random end-effector (Rand EE) birth modes.
@@ IV-A Environment Setup
In both simulation environments, we provide three types of observations: a 2D depth image, a 2D SOI heatmap, and the end-effector position. Based on the observations, the camera and end-effector manipulating the bag are allowed to act simultaneously, but restricted to a maximum of 200 action steps in one episode. Specifically, the actor (i.e. end-effector) can be arbitrarily moved while the active perceiver (i.e., camera) is traversing the half-sphere surface, similar to the CMU’s panoptic massive camera grid  [42] setting. The perceiver agent is controlled by pitch and yaw angles and stares at a constant focus point. To avoid collisions and ensure the simulation stability, we constrain the end-effector action with a bounding box and the camera yaw angle with a range.
@@ IV-B Evaluation Metrics
We use CubeBagClean and CubeBagObst to assess the efficiency of using the proposed action space, in terms of succeeding in finding the cube. In CubeBagClean, we vary the birth modes of the camera and the end-effector to create 4 subtasks where either the camera or the end-effector can be spawned in a fixed or a random place. In CubeBagObst, we consider 2 subtasks with both fixed and random camera birth modes while the end-effector is always initialized in a random position. The fixed camera position is selected in a way that the bag interior is not immediately visible, so non-trivial camera trajectories are required to solve the task. In each task, our reward function encodes a goal of maximizing the visibility of the rigid cube and the opening ring. We use a potential-based reward formulation  [43] to encourage revealing more of the visible cube pixels and the SOI. To demonstrate the effectiveness of DAVS in comparison to the state-of-the-art, we choose a model-free RL algorithm PPO  [44] as the policy search algorithm. We provide quantitative evaluations regarding the finishing episode length and the final reward, and qualitative results illustrating the learned policy behaviors among different methods.
@@ IV-C Baseline Methods
We choose the PPO [44] as our baseline. The 2D depth and heatmap images are concatenated, and embedded as a fix-length feature vector through a convolution neural network architecture from  [45]. The image feature vector and the end-effector location vector are concatenated and used as PPO input. The output camera and end-effector actions thus depend on both current image features and end-effector position.
@@ IV-D3 Generalization to objects with new properties
For each CubeBagClean subtask, we created 100 scenes with random bag dynamics to evaluate the models’ performance, with and without DAVS. As presented in the last two rows of Table I, the IP policy, when utilizing DAVS, not only consistently outperforms its counterpart without DAVS but also demonstrates superior generalizability, underscoring the robustness and effectiveness of our methodology.
Episode Length (normal-$\downarrow$) / Reward (normal-$\uparrow$) / Success Rate(%) (normal-$\uparrow$) Dynamics Method Fix Cam, Fix EE Rand Cam, Fix EE Fix Cam, Rand EE Rand Cam, Rand EE Seen w/o DAVS 200.0 / 0.3 / 0.0 104.4 / 9561.8 / 66.0 164.7 / 3537.0 / 23.0 110.5 / 8958.9 / 49.0 w/ DAVS (ours) 39.0 / 16107.1 / 100.0 64.0 / 13600.7 / 84.0 35.24 / 16483.2 / 95.0 90.1 / 10998.6 / 61.0 Unseen w/o DAVS 200.0 / 0.2 / 0.0 116.4 / 8359.0 / 59.0 152.0 / 4796.1 / 30.0 114.5 / 8557.1 / 46.0 w/ DAVS (ours) 54.4 / 14565.6 / 91.0 84.6 / 11541.6 / 73.0 39.7 / 16035.9 / 94.0 80.7 / 11943.3 / 66.0
TABLE I: Performance across varied Camera and End-Effector Settings
@@ VI Conclusion
In this work, we address deformable object manipulation through Interactive Perception in a system involving an active camera, a robotic end-effector, and a deformable object. We contribute to a novel formulation of the problem with common model assumptions and enable tractable computation through our proposed framework. Then, we perform simulation experiments with different view occlusion situations and demonstrate that our proposed framework with DAVS outperforms the state-of-the-art methods. Finally, we confirm that our method generalizes well to bags featuring previously unseen dynamical properties, as well as their efficacy within an actual real-world interactive perception system.

### v2

@@ I Introduction
Figure 1: An example of Interactive Perception. The perceiver (camera) is moved to a new viewpoint while the actor (end-effector) opens the bag for better perception of in-bag object.
To this end, we propose to address interactive perception for deformable objects in a partially observable Markov Decision Process (POMDP) framework. We focus on bag-like objects and a dual-arm agent setup with an active camera similar to Fig. 1. Our aim is to study the feasibility of solving DOM for IP under the POMDP formulation and whether active vision can facilitate the solution. To leverage the regularity in coupling the perceiver and the actor, our approach proposes to dynamically construct a motion subspace, Dynamic Active Vision Space (DAVS), based on object features for an efficient search. Specifically, we introduce a manifold-with-boundary formulation to characterize a compact camera action subspace to constrain the potential next view. We provide detailed algorithms to calculate the manifold and integrate it into a general action search framework such as reinforcement learning (RL) in solving the POMDP. The simulation results highlight the importance of active vision capability from the controlled camera and the superior performance due to DAVS compared to baseline methods. DAVS shows its strength in generalizing to unseen dynamics and shapes. Finally, we show the proposed method is robust for a successful transfer to the real hardware with a moderate fine-tuning on a dual-arm setup.
@@
  • We conduct extensive studies in both simulation and real-world experiments to demonstrate the effectiveness of our methods in challenging DOM scenarios, We show that our method generalizes well to objects with different dynamical properties and unseen shapes.
@@ II Related Work
Figure 2: [Left] Illustration of the proposed framework — a subspace of camera action is constructed and represented with manifold with boundary, accounting for the coupling with end-effector motion via the structure of interest (SOI) on deformable objects. [Right] Illustration of state action transition of the proposed framework.
@@ III-A Action Space Factorization and Manifold with Boundary
Given the IP task in Fig. 1, we need to propose the camera action $a^{cam}_{t}$ and end-effector action $a^{e}_{t}$ given the observation $o_{t-1}$. A naive solution is to factorize the action space as $\pi(a^{cam}_{t},a^{e}_{t}|o_{t-1})=p(a^{cam}_{t}|o_{t-1})p(a^{e}_{t}|o_{t-1})$ with the assumption of conditional independence, which however results in high dimensionality for action exploration. In this work, we consider that the camera action space should also depend on the previous end-effector action, resulting in $\pi(a^{cam}_{t},a^{e}_{t}|o_{t-1})=p(a^{cam}_{t}|a^{e}_{t-1},o_{t-1})p(a^{e}_{t}|o_{t-1})$. This implies the possibility of sampling $a^{cam}_{t}$ from a subspace depending on $a^{e}_{t-1}$, thereby enabling efficient exploration of the original action space. We use the manifold with boundary $M$ [34] as a tool to describe this subspace and its construction.
The introduction of ${M}$ essentially imposes constraints, and thus structure/regularity, on the original camera action space. To this end, we augment the POMDP $({S},{O},{Z},{T},{R},\gamma)$, such that $z\in{Z}={M}\times{A}^{e}$ is the constrained action space subjecting to ${M}$. The new constrained IP problem can be written as
@@
III-B Manifold with Boundary-based DAVS
@@
Figure 3: Illustration of the process of dynamic active vision space (DAVS) generation. [Left] We extract the SOI points at each time step $t$ and convert them to projected 3D SOI (blue) points through ray tracing, on the original camera action manifold. [Right top] The manifold with boundary based on projected SOI and current camera position by Algorithm 2 [Right bottom] Parameterized exploration space.
@@ 3       Initial parameter setting for active vision ;
5             SOI Modeling $\mathbf{X}(t)\leftarrow f_{SOI}(\boldsymbol{o}_{k-1})$ ;
6             Sample policy end-effector action $\boldsymbol{\alpha}_{k}^{e}\sim\pi(\cdot\mid\boldsymbol{o}_{k-1})$ ;
@@ 5             SOI Modeling $\mathbf{X}(t)\leftarrow f_{SOI}(\boldsymbol{o}_{k-1})$ ;
9             Sample policy camera action $\boldsymbol{\alpha}_{k}^{cam}\sim\pi(\cdot\mid\boldsymbol{o}_{k-1},\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{0},\omega)$ ;
@@ 10             Apply the control camera action $\boldsymbol{\alpha}_{k}^{cam}$ to the environment ;
12             Get next observation $\boldsymbol{o}_{k+1}$ and reward $r_{k}$ from the environment ;
13             Provide the transition tuple $(\boldsymbol{o}_{k-1},\boldsymbol{\alpha}_{k},\boldsymbol{o}_{k},r_{k})$ to the RL algorithm ;
@@ 2 Generate refined active vision space, ${M}^{\prime}(t)$ and its boundary $\partial{M}^{\prime}(t)$ ;
In addition to the centroid, we find two points $P_{1}$ and $P_{2}$ on $\partial{M}^{\prime}(t)$, such that the geodesic path passing through them is locally perpendicular to the one passing through $P_{0}$ and $O_{0}$. As a result, the DAVS is obtained by enclosing the three paths, $\gamma_{\overset{\frown}{P_{0}P_{1}}},\gamma_{\overset{\frown}{P_{1}P_{2}}},\text{and}\gamma_{\overset{\frown}{P_{0}P_{2}}}$. The full process is summarized as Algorithm 2. Additionally, we introduce a $\omega$ to parameterize the exploration space of the camera’s action. As illustrated in Fig. 3, consider the tangent space $T_{P_{0}}M$ at the current camera position $P_{0}$ to a DAVS. Here, the local geodesic paths $\gamma_{\overset{\frown}{P_{0}P_{1}}}$, $\gamma_{\overset{\frown}{P_{0}P_{2}}}$, and $\gamma_{\overset{\frown}{P_{0}O_{0}}}$ are projected onto this tangent space, resulting in vectors $v_{1}$, $v_{2}$, and $v_{0}$. Let us assume that angles $\theta_{1}$ and $\theta_{2}$ are formed between $v_{1}$ and $v_{0}$, and $v_{2}$ and $v_{0}$, respectively. By introducing a parameter $\omega\in[0,1]$, we gain control over the camera action exploration within the range $[\theta_{2}(1-\omega),\theta_{2}+\theta_{1}\omega]$ ($v_{2}$ as zero-degree axis). In our manuscript, we select the action space with $\omega=1$ to provide a broader space for exploring effective policies in complex scenarios. By reducing $\omega$, we can narrow the action sampling space and enforce a preference for alignment with $v_{0}$, which corresponds to the direction along the geodesic path $\gamma_{\overset{\frown}{P_{0}O_{0}}}$. Reducing $\omega$ to 0 causes the camera action policy to strictly follow the geodesic path $\gamma_{\overset{\frown}{P_{0}O_{0}}}$, leading directly to the centroid of DAVS.
@@ III-C Exploiting DAVS in Action Exploration
Figure 4: Left: quantitative evaluation on CubeBagClean. The first row represents the episode length ($\downarrow$) for solving 4 tasks across different combinations of random (Rand) and fixed (Fix) camera (Cam) and end-effector (EE) birth modes. The second row shows the total discounted rewards ($\uparrow$). Right: quantitative evaluation on CubeBagObst subtasks. The first two figures represent the episode length ($\downarrow$) and reward ($\uparrow$) for the scenario with fixed camera (Fix Cam) and random end-effector (Rand EE) birth modes. The third and fourth figures are episode length ($\downarrow$) and reward ($\uparrow$) with random camera (Rand Cam) and random end-effector (Rand EE) birth modes.
@@ IV-A Environment Setup
In both simulation environments, we provide three types of observations: a 2D depth image, a 2D SOI heatmap, and the end-effector position. Based on the observations, the camera and end-effector manipulating the bag are allowed to act simultaneously, but restricted to a maximum of 200 action steps in one episode. Specifically, the end-effector can be arbitrarily moved while the active camera is traversing the half-sphere surface, similar to the CMU’s panoptic massive camera grid  [42] setting. The perceiver agent is controlled by pitch and yaw angles and stares at a constant focus point. To avoid collisions and ensure the simulation stability, we constrain the end-effector action with a bounding box and the camera yaw angle with a range.
@@ IV-B Evaluation Metrics
We use CubeBagClean and CubeBagObst to assess the efficiency of using the proposed action space, in terms of succeeding in finding the cube. In CubeBagClean, we vary the birth modes of the camera and the end-effector to create 4 subtasks where either the camera or the end-effector can be spawned in a fixed or a random place. In CubeBagObst, we consider 2 subtasks with both fixed and random camera birth modes while the end-effector is always initialized in a random position. The fixed camera position is selected in a way that the bag interior is not immediately visible, so non-trivial camera trajectories are required to solve the task. In each task, our reward function encodes a goal of maximizing the visibility of the rigid cube and the opening ring. We use a potential-based reward formulation  [43] to encourage revealing more of the visible cube pixels and the SOI. The reward function is as follows:
$r_{t}=\lambda_{r}\Delta A^{\text{rigid}}_{t}+\lambda_{d}\Delta N^{\text{ring}}_{t}+\lambda_{f}\mathbf{I}(A^{\text{rigid}}_{t},A^{\text{ring}}_{t},t)$ (6)
where $\Delta A^{\text{rigid}}_{t}=A^{\text{rigid}}_{t}-A^{\text{rigid}}_{t-1}$ is the increased visibility of rigid object pixels in the image captured by the active camera. $\lambda_{r}$ is a scale compensation factor depending on the relative distance between the object and the camera. $\Delta N^{\text{ring}}_{t}=N^{\text{ring}}_{t}-N^{\text{ring}}_{t-1}$ is the increased number of visible deformable SOI vertices in the active camera view. $\lambda_{d}$ is a constant scale factor for the deformable object. $\mathbf{I}(\cdot)$ is an indicator function that checks if the IP policy has reached the final goal. Specifically, $\mathbf{I}(\cdot)$ returns 1 only if $A^{\text{rigid}}_{t}$ and $A^{\text{ring}}_{t}$ both reach the preset visibility thresholds before timeout. $\lambda_{f}$ is a time-dependent scale factor, defined as $100\times(T_{\text{max}}-t)$, encouraging the IP policy to finish the task as quickly as possible. To demonstrate the effectiveness of DAVS in comparison to the state-of-the-art, we choose a model-free RL algorithm PPO  [44] as the policy search algorithm. We provide quantitative evaluations regarding the finishing episode length and the final reward, and qualitative results illustrating the learned policy behaviors among different methods.
@@ IV-C Baseline Methods
We choose the PPO [44] and a straightforward Visual Servoing (VS) approach that focuses on maximizing the visibility of the SOI within the camera’s view as our baselines. The 2D depth and heatmap images are concatenated, and embedded as a fix-length feature vector through a convolution neural network architecture from  [45]. The image feature vector and the end-effector location vector are concatenated and used as PPO input. The output camera and end-effector actions thus depend on both current image features and end-effector position.
Episode Length ($\downarrow$) / Reward ($\uparrow$) / Success Rate(%) ($\uparrow$) Dynamics Method Fix Cam, Fix EE Rand Cam, Fix EE Fix Cam, Rand EE Rand Cam, Rand EE Seen w/o DAVS 200.0 / 0.3 / 0.0 104.4 / 9561.8 / 66.0 164.7 / 3537.0 / 23.0 110.5 / 8958.9 / 49.0 w/ DAVS (ours) 39.0 / 16107.1 / 100.0 64.0 / 13600.7 / 84.0 35.24 / 16483.2 / 95.0 90.1 / 10998.6 / 61.0 New Dynamics w/o DAVS 200.0 / 0.2 / 0.0 116.4 / 8359.0 / 59.0 152.0 / 4796.1 / 30.0 114.5 / 8557.1 / 46.0 w/ DAVS (ours) 54.4 / 14565.6 / 91.0 84.6 / 11541.6 / 73.0 39.7 / 16035.9 / 94.0 80.7 / 11943.3 / 66.0 New Shapes w/o DAVS 188.8 / 1125.0 / 11.9 122.5 / 7750.1 / 50.5 95.6 / 10439.2 / 70.5 101.5 / 9859.9 / 58.8 w/ DAVS (ours) 97.9 / 10213.0 / 66.4 98.2 / 10184.4 / 63.1 38.7 / 16141.4 / 93.4 72.7 / 12734.3 / 80.6
TABLE I: Performance across varied Camera and End-Effector Settings
Episode Length ($\downarrow$) / Reward ($\uparrow$) / Success Rate(%) ($\uparrow$) Method CubeBagClean CubeBagObst Total(%) ($\uparrow$) w/o DAVS $137.6$ / $6248.0$ / $36.0$ $87.2$ / 11 539.511539.511\,539.511 539.5 / $65.0$ 50.5 DAVSω=0 32.0 / 16824.8 / 99.0 $143.8$ / $5739.8$ / $36.0$ 67.5 DAVSω=0.5 $59.6$ / 13 950.713950.713\,950.713 950.7 / $82.0$ $90.4$ / 10 760.610760.610\,760.610 760.6 / $59.0$ 71.5 DAVSω=1 $62.7$ / 13 740.913740.913\,740.913 740.9 / $78.0$ 68.3 / 13536.2 / 75.0 76.5 VS $35.4$ / 16 107.716107.716\,107.716 107.7 / $96.0$ $137.5$ / $5348.4$ / $34.0$ 65.0
TABLE II: Performance by Different Approaches
@@ IV-D3 Generalization to objects with new properties
For each CubeBagClean subtask, we created 100 scenes with random bag dynamics to evaluate the models’ performance, with and without DAVS. As presented in Table I, the IP policy, when utilizing DAVS, not only consistently outperforms its counterpart without DAVS but also demonstrates superior generalizability, underscoring the robustness and effectiveness of our methodology.
Furthermore, we investigated the impact of variations in shape, which were not seen during training. Specifically, we created scenes with 36 bags with shapes adapting from [33]. As shown in Table I, the IP policy with DAVS performs better. We observed that certain tasks with unseen shapes were easier to achieve due to two reasons: 1) some new bags had large openings, making it easier to capture the object inside during manipulation; 2) some new bags were more shallow, causing the contained cube to be closer to the camera. Consequently, the rigid cube area appeared relatively larger than in the original setup in the captured image, making reaching the pre-set cube area threshold easier.
IV-D4 Comparison with other baselines
As shown in Table II, with the reduction of $\omega$ to shrink the exploration space for camera actions, the camera is more likely to move directly into the manifold centroid. The PPO policy with DAVS achieves shorter episode lengths, higher total rewards, and higher success rates. While $\boldsymbol{\omega}=0$ significantly improves performance in simpler scenarios (CubeBagClean), higher values of $\omega$ (DAVSω=0.5 and DAVSω=1 ) are crucial for handling more complex environments (CubeBagObst). The Visual Servoing method (VS) performs better than DAVSω=1 in some aspects but is outperformed by DAVS with $\omega$ settings on CubeBagObst. In general, DAVSω=1 achieves the best overall balance, making it the most robust approach among the tested methods. We also examined the failure cases in the simulation experiments. In our Pybullet simulation setup, we executed the movements of the camera and the actor simultaneously. Most failures occurred when the simulated gripper moved too quickly, causing overstretching of the deformable handles. Additionally, the relatively slow movement of the camera made it more difficult to locate the object, especially in scenes involving obstacles. Furthermore, under the challenging generalization test, the IP method failed particularly when the stiffness differed significantly or with a small bag opening.
@@ VI Conclusion
In this work, we address deformable object manipulation through Interactive Perception in a system involving an active camera, a robotic end-effector, and a deformable object. We contribute to a novel formulation of the problem with common model assumptions and enable tractable computation through our proposed framework. Then, we perform simulation experiments with different view occlusion situations and demonstrate that our proposed framework with DAVS outperforms the state-of-the-art methods. Finally, we confirm that our method generalizes well to bags featuring previously unseen dynamical properties, as well as their efficacy within an actual real-world interactive perception system. The challenges presented in this work lead to the future directions of improving hand-eye coordination, such as explicitly incorporating dynamics prediction within the IP framework, and more effective parameterized DAVS constructions with task dependency for deformable object manipulation.